OpenStudy (anonymous):

Which of the following is not a step used when graphing the equation x + 3y = 3 using the x- and y- intercepts method?

OpenStudy (anonymous):

it's a linear equation , so what do you want exactly ? , it intercept y-axis in at 3

OpenStudy (johnweldon1993):

Actually it intercepts the y-axis at 1 to find the y-intercept...make x = 0 x + 3y = 3 0 + 3y = 3 ------- 3 3 y = 1 This is your y-intercept to find the x-intercept...make y = 0 x + 3y = 3 x + 0 = 3 x = 3 This is your x-intercept

OpenStudy (anonymous):

sry u rigth i forgot to dvide by 3 :)

OpenStudy (anonymous):

Plot the y-intercept (0, 1). Plot the intercept (3, 0). Count down 1 and right 3 from the y-intercept. Draw a line through the two points. those are my options and I have to choose whats not the choice so wouldn't it be the third choice?

OpenStudy (johnweldon1993):

If it is ONLY using the x and y intercept method...then you need not worry about the slope....so yes the third one :)

OpenStudy (anonymous):

Okay Thank You (:

OpenStudy (anonymous):

can you help me with another one?